# Normal forms for ordinary differential operators, I

### v3

@@ 3 Normal forms for commuting operators
where $I=(f_{1},\ldots,f_{k})$ is a prime ideal and the order of $w_{1}$ is coprime with the order of $w_{2}$, and the images of $w_{i}$ under the Krichever map (after some choice of $\pi$, $\hat{\phi}$) are monic (cf. remark 3.7).
Then there exist normalised normal forms of rank one $B^{\prime}\simeq{{\cal O}}_{C}(C_{0})$ with respect to the ordered set of generators $P_{1}^{\prime}=\partial^{q},\ldots,P_{m}^{\prime}$, where $\mathop{\rm\bf ord}(P_{i}^{\prime})=\mathop{\rm ord}\nolimits(w_{i})$ for all $i\geq 1$, and there is a one to one correspondence between closed points of the affine algebraic set $X_{[{{\cal O}}_{C}(C_{0})]}$ and isomorphism classes of torsion free rank one sheaves ${{\cal F}}$ on $C$ with vanishing cohomologies $H^{0}(C,{{\cal F}})=H^{1}(C,{{\cal F}})=0$.
@@
Let ${{\cal F}}$ be a torsion free rank one sheaf on $C$ with vanishing cohomologies. By theorem 3.1 the triple $(C,p,{{\cal F}})$ corresponds to uniquely defined normalised [^9] Each equivalence class $[B]$ contains a normalised representative, which determines the whole class uniquely. For example, we can choose a monic differential operator of minimal positive order in $B$, and normalise it (conjugating by an appropriate function $f\in\hat{R}$). commutative elliptic subring $B\subset D_{1}$ up to a scale transform. Besides, the generators $w_{i}$ corresponds to formally elliptic differential operators $P_{1},\ldots,P_{m}$ of the same orders, and there exists a scale transform that makes them monic. By proposition 2.3 and lemma 3.4 there exists a uniquely determined Schur operator $S\in\hat{D}_{1}^{sym}$ such that $B^{\prime}:=S^{-1}BS\in C(\partial^{q})$ and $P_{2}^{\prime}=S^{-1}P_{2}S$ is a normalised normal form of $P_{2}$ with respect to $P_{1}$, i.e. $B^{\prime}$ is a normalised normal form w.r.t. $P_{1}^{\prime},\ldots,P_{m}^{\prime}$ which have the same orders as $w_{i}$ or $P_{i}$. Note that the scale transform is compatible with conjugation by $S$ and that the coefficients of $\hat{\Phi}(P_{i}^{\prime})$ are invariant under any scale transform of $P_{i}^{\prime}$ for all $i$. Thus, ${{\cal F}}$ determines a closed point of $X_{[{{\cal O}}_{C}(C_{0})]}$.
Vice versa, any closed point of $X_{[{{\cal O}}_{C}(C_{0})]}$ determines a normalised normal form $B^{\prime}$ w.r.t. some $P_{1}^{\prime},\ldots,P_{m}^{\prime}$ which have the same orders as $w_{i}$. By corollary 3.1 and remark 3.5 there exists an operator $S\in E_{q}$ such that $A^{\prime}:=S^{-1}\hat{\Phi}(B^{\prime})S\in K((\tilde{D}^{-1}))$. Note that $W^{\prime\prime}:=F^{\prime\prime}\cdot S$ has support equal to $F^{\prime\prime}$, because $S$ is a monic invertible operator of order zero. So, $(W^{\prime\prime},A^{\prime})$ form an embedded Schur pair of rank one [^10] Since $\hat{\Phi}(C(\partial^{p}))$ consists of Laurent polynomials whose coefficients at negative powers of $\tilde{D}$ are not constants (see remark 3.4), and since $S$ is monic and $A^{\prime}\subset K((\tilde{D}^{-1}))$, it follows that $A^{\prime}\cap K[[\tilde{D}^{-1}]]=K$.. By theorems 3.2 and 3.1 this Schur pair determines a normalised commutative subring $B\subset D_{1}$ of rank one and a torsion free sheaf ${{\cal F}}\simeq\mathop{\rm Proj}\tilde{W^{\prime\prime}}$ of rank one with vanishing cohomologies.
@@
Let $L=\partial^{2}+u$, $P=4\partial^{3}+6u\partial+3u^{\prime}$, where $u(x)=\sum_{k=0}^{\infty}\frac{1}{k!}u_{k}\cdot x^{k}$, be ODOs of orders 2 and 3. Then $[L,P]=0$ iff $06uu^{\prime}+u^{\prime\prime\prime}=$ 0 (see [41]). It’s easy to see that the coefficients $u_{k}$, $k\geq 3$ are uniquely determined by this equation for any choice of free parameters $u_{0},u_{1},u_{2}$. The spectral curve of these operators is given by
@@
$\tilde{P}=\psi\circ\hat{\Phi}(P)=\left(\begin{array}[]{cc}u_{1}&2\left(2\tilde{D}^{2}+u_{0}\right)\\ 4\tilde{D}^{4}-2\tilde{D}^{2}u_{0}-2u_{0}^{2}-u_{2}&-u_{1}\\ \end{array}\right)$
@@
On the other hand, if we have two operators $L_{2},L_{3}\in D_{1}$ satisfying the equation $L_{3}^{2}=4L_{2}^{3}-g_{2}L_{2}-g_{3}$, where $g_{2},g_{3}$ are two constant coefficients in $K$ (may assume $L_{2}$ is normalized), then apparently $[L_{2},L_{3}]=0$ (cf. [7]).
Consider the normal form $L_{3}^{\prime}$ of $L_{3}$ with respect to $L_{2}$. We fix $k=2,\xi=-1$ in this example, according to Lemma 2.4, $L_{3}^{\prime}$ should be in the form of (within $Hcpc(2)$):
@@
The calculations of $L_{3}^{\prime}$ will be basically separated into the following parts:
@@
    Substitute the expression of $(L_{2},L_{3})$ into equation $L_{3}^{2}=4L_{2}^{3}-g_{2}L_{2}-g_{3}$, we get the relation that

### v4

@@ 3 Normal forms for commuting operators
where $I=(f_{1},\ldots,f_{k})$ is a prime ideal and the order of $w_{1}$ is coprime with the order of $w_{2}$, and the images of $w_{i}$ under the Krichever map (after some choice of $\pi$) are monic (cf. remark 3.7).
[…]
@@
[…]
Vice versa, any closed point of $X_{[{{\cal O}}_{C}(C_{0})]}$ determines a normalised normal form $B^{\prime}$ w.r.t. some $P_{1}^{\prime},\ldots,P_{m}^{\prime}$ which have the same orders as $w_{i}$. By corollary 3.1 and remark 3.5 there exists an operator $S\in E_{q}$ such that $A^{\prime}:=S^{-1}\hat{\Phi}(B^{\prime})S\in K((\tilde{D}^{-1}))$. Note that $W^{\prime\prime}:=F^{\prime\prime}\cdot S$ has support equal to $F^{\prime\prime}$, because $S$ is a monic invertible operator of order zero. So, $(W^{\prime\prime},A^{\prime})$ form an embedded Schur pair of rank one [^15] Since $\hat{\Phi}(C(\partial^{q}))$ consists of Laurent polynomials whose coefficients at negative powers of $\tilde{D}$ are not constants (see remark 3.4), and since $S$ is monic and $A^{\prime}\subset K((\tilde{D}^{-1}))$, it follows that $A^{\prime}\cap K[[\tilde{D}^{-1}]]=K$.. By theorems 3.2 and 3.1 this Schur pair determines a normalised commutative subring $B\subset D_{1}$ of rank one and a torsion free sheaf ${{\cal F}}\simeq\mathop{\rm Proj}\tilde{W^{\prime\prime}}$ of rank one with vanishing cohomologies.
@@
Let $L=\partial^{2}+u$, $P=4\partial^{3}+6u\partial+3u^{\prime}$, where $u(x)=\sum_{k=0}^{\infty}\frac{1}{k!}u_{k}\cdot x^{k}$, be ODOs of orders 2 and 3. Then $[L,P]=0$ iff $6uu^{\prime}+u^{\prime\prime\prime}=0$ (see [44]). It’s easy to see that the coefficients $u_{k}$, $k\geq 3$ are uniquely determined by this equation for any choice of free parameters $u_{0},u_{1},u_{2}$. The spectral curve of these operators is given by
@@
$\tilde{P}^{\prime}=\psi\circ\hat{\Phi}(P^{\prime})=\left(\begin{array}[]{cc}u_{1}&2\left(2\tilde{D}^{2}+u_{0}\right)\\ 4\tilde{D}^{4}-2\tilde{D}^{2}u_{0}-2u_{0}^{2}-u_{2}&-u_{1}\\ \end{array}\right)$
@@
On the other hand, if we have two operators $L_{2},L_{3}\in D_{1}$ satisfying the equation $L_{3}^{2}=4L_{2}^{3}-g_{2}L_{2}-g_{3}$, where $g_{2},g_{3}$ are two constant coefficients in $K$ (may assume $L_{2}$ is normalized), then apparently $L_{2},L_{3}]=0$ (cf. [7]).
Consider the normal form $L_{3}^{\prime}$ of $L_{3}$ with respect to $L_{2}$. We fix $k=2,\xi=-1$ in this example, according to Lemma 2.5, $L_{3}^{\prime}$ should be in the form of (within $E_{2}$):
@@
The calculations of $L_{3}^{\prime}$ can be separated into the following parts:
@@
    Substitute the expression of $(L_{2},L_{3})$ into equation $L_{3}^{2}=4L_{2}^{3}-g_{2}L_{2}-g_{3}$, we get the relation that [^17] We can substitute $\hat{\Phi}(L_{3})$ and $\tilde{D}^{2}$ instead, in this case calculations are easy in both cases.